Please help, I'm so confused!!!! This is due wednesday night!!! i'm gonna fail :((((pH of Buffer Solutions Procedure: Acetic Acid-Sodium Acetate Buffer (pKa acetic acid = 4.75) Weigh about 3.5 g of Na2C2H302 3Week 9 pH of Buffer Solutions Data Show all work Measured pH (with pH meter) of original buffer solution 4.80 Acetic Acid-SodIf acid is added to a buffer would the pH increase or decrease? pH of buffer + NaOH 5.21 Calculated pH of buffer + NAOH: Repe
Answer #1

Acetic acid-sodium acetate buffer:
Measured pH of original buffer solution = 4.80
1. moles of sodium acetate:
moles, nNaOAc = mass of sodium acetate /molar mass of sodium acetate
nNaOAc = 3.51/136 = 0.0258 moles
Calculated pH of original buffer solution:
Total volume of solution: 55.6 mL water + 8.8 mL 3M acetic acid (HA) = 64.4 mL
2. Calculate M [A-]: Molarity = moles/volume
[A-] = 0.0258/0.0644 = 0.40 M
3. Calculate M [HA]:

using the dilution relation as M1V1 = M2V2
M1 =stock concnetration of HA, V1 = volume of HA in solution, V2 = solution volume
thus [HA] in solution, M2 = M1V1 /V2
[HA] = 8.8*3/64.4 = 0.41 M
4. pKa of acetic acid: 4.74
5. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 4.74 + log(0.40/0.41)
pH = 4.74 - 9.9*10-3
pH = 4.73
Addition of Acid to buffer

Measured pH of buffer + HCl = 4.24
Calculated pH of buffer + HCl solution:
Total volume of solution: 32.2 + 1 mL 6M HCl = 33.2 mL
1. calculate M HCl added: 1*6/33.2 = 0.18 M HCl
2. new[A-] = 0.40 M [A-] - 0.18 M [HCl] = 0.22 M
3. new [HA] = 0.41 M [HA] + 0.18 M [HCl] = 0.59 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 4.74 + log(0.22/0.59)
pH = 4.74 - 0.43
pH = 4.31
Upon addition of acid, pH of decreases.
Addition of NaOH to buffer

pH of buffer + NaOH = 5.21
Calculated pH of buffer + NaOH solution:
Total volume of solution: 32.2 + 1 mL 6M NaOH = 33.2 mL
1. calculate M NaOH added: 1*6/33.2 = 0.18 M NaOH
2. new[A-] = 0.40 M [A-] + 0.18 M [NaOH] = 0.58 M
3. new [HA] = 0.41 M [HA] - 0.18 M [NaOH] = 0.23 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 4.74 + log(0.58/0.23)
pH = 4.74 + 0.40
pH = 5.14
Upon addition of base, pH of increases.

Carbonic acid-sodium hydrogen carbonate buffer system

System 1:1

stock solution [HA] = [A-] = 0.1 M

Calculated pH of original buffer solution:
Total volume of solution: 32.0 mL HA + 32.0 mL 0.1M A- = 64.0 mL
1. Calculate M [A-]: using the dilution relation as M1V1 = M2V2
M1 = stock concentration of A-, V1 = volume of A- in solution, V2 = solution volume
thus [A-] in solution = M1V1/V2
[A-] = 0.1*32.0/64.0 = 0.05 M
2. Calculate M [HA]:
for 1:1 system [HA] = [A-] = 0.05 M
3. pKa of carbonic acid: 6.37
5. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.05/0.05)
pH = 6.37 + 0
pH = 6.37

Calculated pH of buffer + HCl solution:
Total volume of solution: 32.0 + 0.1 mL 6M HCl = 32.1 mL
1. calculate M HCl added: 0.1*6/32.1 = 0.0187 M HCl
2. new[A-] = 0.05 M [A-] - 0.0187 M [HCl] = 0.0313 M
3. new [HA] = 0.05 M [HA] + 0.0187 M [HCl] = 0.0687 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.0313/0.0687)
pH = 6.37 - 0.34
pH = 6.03
Upon addition of acid, pH of decreases.

Calculated pH of buffer + NaOH solution:
Total volume of solution: 32.0 + 0.1 mL 6M NaOH = 32.1 mL
1. calculate M NaOH added: 0.1*6/32.1 = 0.0187 M NaOH
2. new[A-] = 0.05 M [A-] + 0.0187 M [NaOH] = 0.0687 M
3. new [HA] =0.05 M [HA] - 0.0187 M [NaOH] = 0.0313 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.0687/0.0313)
pH = 6.37 + 0.34
pH = 6.71
Upon addition of base, pH of buffer increases.

System 1:10 [A-] = 10[HA]

stock solution [HA] = [A-] = 0.1 M

Calculated pH of original buffer solution:
Total volume of solution: 6.4 mL HA + 64 mL 0.1M A- = 70.4 mL
1. Calculate M [A-]: using the dilution relation as M1V1 = M2V2
M1 = stock concentration of A-, V1 = volume of A- in solution, V2 = solution volume
thus [A-] in solution = M1V1/V2
[A-] = 0.1*64/70.4 = 0.091 M
2. Calculate M [HA]:
for 1:10 system [HA] = 0.1* [A-] = 0.0091 M
3. pKa of carbonic acid: 6.37
5. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.091/0.0091)
pH = 6.37 + 1
pH = 7.37

Calculated pH of buffer + HCl solution:
Total volume of solution: 35.2 + 0.1 mL 6M HCl = 35.3 mL
1. calculate M HCl added: 0.1*6/35.3 = 0.017 M HCl
2. new[A-] = 0.091 M [A-] - 0.017 M [HCl] = 0.074 M
3. new [HA] = 0.0091 M [HA] + 0.017 M [HCl] = 0.108 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.074/0.108)
pH = 6.37 - 0.16
pH = 6.21
Upon addition of acid, pH of decreases.

Calculated pH of buffer + NaOH solution:
Total volume of solution: 35.2 + 0.1 mL 6M NaOH = 35.3 mL
1. calculate M NaOH added: 0.1*6/35.3 = 0.017 M NaOH
2. new[A-] = 0.091 M [A-] + 0.017 M [NaOH] = 0.108 M
3. new [HA] =0.0091 M [HA] - 0.017 M [NaOH] = 0.074 M
4. pH = pKa+log ([A-]/[HA])
substituting the values we get:
pH = 6.37 + log(0.108/0.074)
pH = 6.37 + 0.16
pH = 6.53
Upon addition of base, pH of increases.

Water as buffer:

[H2O] = 55.55 M

In water, [OH-] = [H+]10-7 M

pH = -log [H+]

pH = 7

Addition of 1 mL of 6 M HCl

Volume of solution = 32 mL water + 1 mL of 6 M HCl = 33 mL

[H+] due to 1 mL of 6 M HCl = 1*6/33 = 0.182 M

total [H+] = 10-7 M + 0.182 M = 0.182 M

pH = -log (0.182)

pH = 0.74

pH of water decreases upon addition of NaOH.

Addition of 1 mL of 6 M NaOH

Volume of solution = 32 mL water + 1 mL of 6 M NaOH = 33 mL

[OH-] due to 1 mL of 6 M naOH = 1*6/33 = 0.182 M

total [OH-] = 10-7 M + 0.182 M = 0.182 M

pOH = -log (0.182)

pOH = 0.74

using the formula to relate pH and pOH

pH = 14 - pOH

pH = 14 -0.74

pH = 13.26

pH of water increases upon addition of acid.